How do you solve z+81=9z-7
Answer by vertciel(183)

z + 81 = 9z - 7
+ 7 + 81 = 9z - z
9z - z = 81 + 7
8z = 88
z = 11

I think you made some other posts on algebraic expressions as well. Try to do them yourself since I showed you how to do this one.